OpenStudy (anonymous):

pleaz help me... to convert women's shoe sizes in the us to shoe sizes in italy, use the function f(s)=2s+24. to convert shoes sizes in italy to shoe sizes in the Great britain, use the function g(s)=1/2s n 14. what is the function for converting women's shoe sizes in the us to shoe sizes in great britain

OpenStudy (anonymous):

f(g(s)) so = 2*(1/2s n 14)+24, dont know why there is an n but you see what i mean

OpenStudy (anonymous):

well the answer choices r a. f(g)(s)= s+7 b. (f+g)(s)=21/2s+10 c. g(f)(s)=s-2 d. fg(s)=s^2-16s-336

OpenStudy (anonymous):

can you rewrite g(s)

OpenStudy (anonymous):

sure g(s)=1/2s n14

OpenStudy (anonymous):

why is there an n. it cant be a variable - none in solutions. i need to see the question.

OpenStudy (anonymous):

well thats how it says. to convert women;s shoe sizes in the us to shoe sizes in italy, use the function f(s)=2s+24. to convert shoe sizes in italy to shoe sizes in G.B., use the function g(s)=1/2s n 14. what is a function for converting women's shoe sizes in Great Britain?

OpenStudy (anonymous):

1/2s n 14 is my whole point, what is the n operator

OpenStudy (anonymous):

i dont get it myself, i've no clue. :(

OpenStudy (anonymous):

you should take a pic of the problem.

OpenStudy (anonymous):

dont have a camera :(

OpenStudy (anonymous):

i typed it word 4 word though.